# Polylogarithmic motivic Chabauty–Kim for $\mathbb{P}^{1}\setminus\{0,1,\infty\}$: the geometric step via resultants

### v2

@@ Acknowledgements
The authors thank David Corwin and Ishai Dan-Cohen for their support and encouragement throughout the project. They also thank Martin Lüdtke for numerous helpful conversations, and both Martin Lüdtke and Alexander Betts for their comments and corrections on a preliminary version of this paper. DJ and SZ were supported by the Israel Science Foundation (grants No. 726/17 and No. 621/21 respectively, PI: Ishai Dan-Cohen) while at Ben-Gurion University of the Negev. DTBGL was supported by an Emily Erskine Endowment Fund Post-Doctoral Researcher Fellowship and the Israel Science Foundation (grant No. 2301/20, PI: Ari Shnidman) while at the Hebrew University of Jerusalem, and by an Edixhoven Post-Doctoral Fellowship while at Leiden University. FMS and AW were supported by the Israel Science Foundation (grant No. 1963/20, PI: Daniel Disegni) and by the US-Israel Binational Science Foundation (grant No. 2018250, PI: Daniel Disegni) while at Ben-Gurion University of the Negev.
@@ 2.1. Mixed Tate Motives
where, for each integer $n$, the quotient $\operatorname{gr}_{n}^{W}(Y)=W_{n}(Y)/W_{n-1}(Y)$ is an element of the sub-tensor category $\operatorname{MT}(\mathbb{Z}_{S})_{n}$ of $\operatorname{MT}(\mathbb{Z}_{S})$ spanned by $\mathbb{Q}(-n/2)$ (we adopt the convention that $\mathbb{Q}(-n/2)=0$ if $n$ is odd). The above filtration is called the weight filtration.
@@ 2.2. The pro-unipotent mixed Tate fundamental group
gives a non-trivial map from $G_{S}$ to $\operatorname{GL}(\mathbb{Q})$, and thus a surjection $G_{S}\twoheadrightarrow\mathbb{G}_{m}$.
@@
of affine algebraic groups over $\mathbb{Q}$. This short exact sequence splits: letting $\lambda\in\mathbb{G}_{m}$ act on $\omega_{d}(X)$ by multiplication by $\lambda^{d}$ gives a $\mathbb{G}_{m}$-action (i.e. a grading) on $\omega(X)$, and hence a splitting $\mathbb{G}_{m}\to G_{S}$ (known as the Levy splitting of $G_{S}$). Thus, there is an isomorphism
@@
We may describe $U_{S}$ in terms of the extension groups $\operatorname{Ext}^{1}(\mathbb{Q}(0),\mathbb{Q}(n))$ in $\operatorname{MT}(\mathbb{Z}_{S})$ (the $\operatorname{Ext}^{2}$ are trivial in $\operatorname{MT}(\mathbb{Z}_{S})$ [delignegoncharov, Prop. 1.9]), as we will now explain. By (2.1), we have
@@
When $n=1$, we have $K_{1}(\mathbb{Z}_{S})\otimes\mathbb{Q}=\mathbb{Z}_{S}^{\times}\otimes\mathbb{Q}$, which has dimension $|S|$. When $n\geq 2$, we have $K_{2n-1}(\mathbb{Z}_{S})\otimes\mathbb{Q}=K_{2n-1}(\mathbb{Q})\otimes\mathbb{Q}$, which has dimension 00 for $n$ even and dimension $1$ for $n$ odd [borel].
By the formalism of pro-unipotent groups, it follows that $\mathrm{Lie}(U_{S})$ is a free, negatively graded, pro-nilpotent Lie algebra, which is finite-dimensional in each degree (the grading is negative because of Remark 2.2). Moreover, the functor $\omega$ induces an equivalence between $\operatorname{MT}(\mathbb{Z}_{S})$ and the category of graded finite-dimensional $\mathbb{Q}$-vector spaces with an action of $\operatorname{Lie}(U_{S})$ compatible with the gradings [delignegoncharov, Prop. 2.2]. By [delignegoncharov, A.15], a system of homogeneous free generators of $\operatorname{Lie}(U_{S})$ is obtained by (non-canonically) lifting a basis of each
@@
Hence, the pro-unipotent group $U_{S}$ is non-canonically isomorphic to the free pro-unipotent group (in the sense of [CDC1, §2.1.3]) on the graded vector space $\bigoplus_{n\in\mathbb{Z}}\operatorname{Ext}^{1}(\mathbb{Q}(0),\mathbb{Q}(n))^{\vee}$ with $\operatorname{Ext}^{1}(\mathbb{Q}(0),\mathbb{Q}(n))^{\vee}$ placed in degree $-n$.
@@
Fix once and for all a system of free generators of $\operatorname{Lie}(U_{S})$ given by $\{\tau_{\ell}\}_{\ell\in S}$ in degree (i.e., half-weight) $-1$, and $\sigma_{2n+1}$ in degree $-2n-1$, for each $n\geq 1$. We denote by $\Sigma$ this ordered set of free generators.
@@ 2.3. The unipotent de Rham fundamental group
Fix a point $x\in\mathbb{Q}\setminus\{0,1\}$ or a tangential vector at $0,1$ or $\infty$. The unipotent de Rham fundamental group of $X$ at $x$, $\pi^{\mathrm{un},\operatorname{dR}}_{1}(X;x)$, is the Tannakian fundamental group of the neutral $\mathbb{Q}$-Tannakian category of algebraic vector bundles with nilpotent connection on $X_{\mathbb{Q}}$, with fibre functor given by taking the fibre at $x$. In the case that $x=\vec{1}_{0}$ is the tangential vector $1$ at the point 00, we set
@@
The group $\Pi$ is an affine, pro-unipotent group over $\mathbb{Q}$. Its coordinate ring $\mathcal{O}(\Pi)$ is the de Rham realisation of a commutative Hopf algebra in the category of $\mathrm{ind}$-objects of $\operatorname{MT}(\mathbb{Z})$, the category of unramified mixed Tate motives over $\mathbb{Q}$ [delignegoncharov, §4, §5]. The de Rham realisation functor of $\operatorname{MT}(\mathbb{Z}_{S})$ is canonically isomorphic to $\omega$ [delignegoncharov, §2.9]. In particular, the group $\Pi$ carries an action of $G_{S}$.
Let $V$ denote the graded vector space consisting of $H_{1}^{\operatorname{dR}}(X_{\mathbb{Q}})$ in degree $-1$ and 00 in all other degrees. Then we can view $\Pi$ as the free pro-unipotent group on $V$, in the following sense. Take the completed graded tensor algebra $TV$ on $V$ and put the unique coproduct on it such that all elements of $V$ are primitive. Denote by $\mathfrak{n}$ the subspace of $TV$ consisting of primitive elements. Then $\mathfrak{n}$ has the structure of a strictly negatively graded pro-nilpotent Lie algebra, so it is the Lie algebra of a graded pro-unipotent group $U(V)$. The statement that $\Pi$ is the free pro-unipotent group on $V$ means that $\Pi=U(V)$. In particular, the coordinate ring $\mathcal{O}(\Pi)$ is the Hopf dual algebra of the universal enveloping algebra $\mathcal{U}\mathfrak{n}$. This enveloping algebra is $TV$ with the unique coproduct for which all elements of $H_{1}^{\operatorname{dR}}(X_{\mathbb{Q}})$ are primitive.
This description allows us to apply the machinery of [CDC1, §2.1.3]. Take $\{e_{0},e_{1}\}\subseteq TV$ to be the basis dual to the standard basis $\{dz/z,dz/(1-z)\}$ of $H^{1}_{\operatorname{dR}}(X_{\mathbb{Q}})$, assign to these basis elements the degree $-1$ and declare them to be primitive. Then $TV$ is generated by words $w$ in the basis elements $e_{0}$ and $e_{1}$. Let $\operatorname{Li}^{\mathfrak{u}}_{w}$ denote the element of $\mathcal{O}(\Pi)$ that is dual to the word $w$. Then $\mathcal{O}(\Pi)$ is generated by the elements $\{\operatorname{Li}^{\mathfrak{u}}_{w}\}_{w}$. The deconcatenation coproduct is given by
@@ 3.1. The polylogarithmic quotient
Recall from Section 2.3 that $\Pi$ is the free pro-unipotent group on the graded vector space $V$, which is $H_{1}^{\operatorname{dR}}(X_{\mathbb{Q}})=\mathrm{Span}_{\mathbb{Q}}\{e_{0},e_{1}\}$ in degree $-1$ and zero elsewhere. The specific words of the form
@@ 3.2. The Selmer scheme
$\operatorname{Sel}_{S}^{\operatorname{PL}}(\mathbb{Q})=H^{1}(G_{S},\Pi^{\operatorname{PL}}):=\{\Pi^{\operatorname{PL}}\text{-torsors over }\mathbb{Q}\text{ with a compatible action of }G_{S}\}.$
@@ 3.3. The Chabauty–Kim diagram
    $\varphi^{\#}_{b}\colon\mathcal{O}(\Pi)\to\mathbb{Q}_{p},\qquad\operatorname{Li}^{\mathfrak{u}}_{w}\mapsto\int_{\vec{1}_{0}}^{b}\omega_{w},$
@@ 3.4. The Chabauty–Kim locus
Recall from Definition 3.3 that $\mathcal{I}^{\operatorname{PL}}_{S}=\{f\in\mathcal{O}((\Pi^{\operatorname{PL}})_{K})\mid f(\operatorname{Im}(\operatorname{ev}_{K}))=0\}$. Given an integer $d\geq 1$, we define
@@
Moreover, the set $X(\mathbb{Z}_{p})^{\mathrm{PL}}_{S,d}$ is finite for sufficiently large $d$ (this follows for instance from Proposition 5.2 below). It is conjectured [CDC1, Conj. 2.32] that
(3.4) $X(\mathbb{Z}_{S})=(X(\mathbb{Z}_{p})^{\mathrm{PL}}_{S,d})^{S_{3}},\qquad\text{for }d\gg 0.$
@@
where $X(\mathbb{Z}_{p})^{\rm Kim}_{S,d}$ denotes Kim’s $p$-adic Chabauty locus in depth $d$ (see [CDC1, Rem. 2.28]). From these inclusions, it is clear that conjecture (3.4) implies Kim’s conjecture [kim]. One insight of [CDC1] is that is suffices to use single polylogarithms to deal with the $S$-unit equation, as opposed to requiring multiple polylogarithms as in Kim’s original approach. Dan-Cohen and Corwin verified the conjectural equality (3.4) for $S=\{3\}$, $p=5,7$, and $d=4$ [CDC1, Thm. 5.5].
@@
  • Arithmetic step in depth $d$: the task of determining the $p$-adic Chabauty–Kim functions $f|_{X(\mathbb{Z}_{p})}$ defined by (3.2) for the functions $f$ found in the geometric step. This requires the calculation of the image of $f$ under the $p$-adic period map $\operatorname{per}_{p}$. The resulting functions $f|_{X(\mathbb{Z}_{p})}$ are guaranteed to be non-trivial if one assumes the $p$-adic Period Conjecture [CDC1, Conj. 2.25]. This step depends on the prime $p$ and the specific primes contained in $S$.
@@ 4. The geometric step in coordinates
where we recall that the ideal $\mathcal{I}^{\operatorname{PL}}_{S}=\ker(\operatorname{ev}_{K}^{\#})$ is defined in Definition 3.3. In this section we describe, following [CDC1], a (non-canonical) affine model for the Selmer scheme $\operatorname{Sel}_{S}^{\operatorname{PL}}$ (Definition 3.2) and give an explicit formula for the map $\operatorname{ev}$ in the chosen coordinates.
@@ 4.1. Coordinates on Selmer schemes
Recall from Definition 2.5 the fixed system of free generators $\Sigma=\{\tau_{\ell},\sigma_{2n+1}\}$ of $\operatorname{Lie}(U_{S})$, with their specified degrees. Each word $w$ in these generators corresponds to a dual element in $\mathcal{O}(U_{S})$, denoted $f_{w}$ (the abstract shuffle basis).
@@ 4.2. The evaluation map in coordinates
$\log^{\mathfrak{u}}\mapsto\sum_{\tau\in\Sigma_{-1}}f_{\tau}\Phi_{e_{0}}^{\tau}\qquad\text{ and }\qquad\operatorname{Li}_{k}^{\mathfrak{u}}\mapsto\sum f_{\sigma\tau_{(1)}\ldots\tau_{(r)}}\Phi^{\sigma}_{e_{1}(e_{0})^{s-1}}\Phi_{e_{0}}^{\tau_{(1)}}\ldots\Phi_{e_{0}}^{\tau_{(r)}},$
where the latter sum is taken over $\tau_{(1)},\ldots,\tau_{(r)}\in\Sigma_{-1}$ and $\sigma\in\Sigma_{-s}$ such that $r+s=k$ and $1\leq s\leq n$.

### v3

@@ Acknowledgements
The authors thank David Corwin and Ishai Dan-Cohen for their support and encouragement throughout the project. They also thank Martin Lüdtke for numerous helpful conversations, and both Martin Lüdtke and Alexander Betts for their comments and corrections on a preliminary version of this paper, as well as the anonymous referees. DJ and SZ were supported by the Israel Science Foundation (grants No. 726/17 and No. 621/21 respectively, PI: Ishai Dan-Cohen) while at Ben-Gurion University of the Negev. DTBGL was supported by an Emily Erskine Endowment Fund Post-Doctoral Researcher Fellowship and the Israel Science Foundation (grant No. 2301/20, PI: Ari Shnidman) while at the Hebrew University of Jerusalem, and by an Edixhoven Post-Doctoral Fellowship while at Leiden University. FMS and AW were supported by the Israel Science Foundation (grant No. 1963/20, PI: Daniel Disegni) and by the US-Israel Binational Science Foundation (grant No. 2018250, PI: Daniel Disegni) while at Ben-Gurion University of the Negev.
@@ 2.1. Mixed Tate Motives
where, for each integer $n$, the quotient $\operatorname{gr}_{n}^{W}(Y)=W_{n}(Y)/W_{n-1}(Y)$ is isomorphic to a direct sum of a finite number of copies of $\mathbb{Q}(-n/2)$ (we adopt the convention that $\mathbb{Q}(-n/2)=0$ if $n$ is odd). The above filtration is called the weight filtration.
@@ 2.2. The pro-unipotent mixed Tate fundamental group
gives a non-trivial map from $G_{S}$ to $\operatorname{GL}_{1,\mathbb{Q}}$, and thus a surjection $G_{S}\twoheadrightarrow\mathbb{G}_{m}$.
@@
of affine algebraic groups over $\mathbb{Q}$. This short exact sequence splits: letting $\lambda\in\mathbb{G}_{m}$ act on $\omega_{d}(X)$ by multiplication by $\lambda^{d}$ gives a $\mathbb{G}_{m}$-action (i.e., a grading) on $\omega(X)$, and hence a splitting $\mathbb{G}_{m}\to G_{S}$ (known as the Levy splitting of $G_{S}$). Thus, there is an isomorphism
@@
We may describe $U_{S}$ in terms of the extension groups $\operatorname{Ext}^{1}(\mathbb{Q}(0),\mathbb{Q}(n))$ in $\operatorname{MT}(\mathbb{Z}_{S})$ (the $\operatorname{Ext}^{2}$ are trivial in $\operatorname{MT}(\mathbb{Z}_{S})$ [delignegoncharov, Prop. 1.9]), as we will now explain.  [^3] To avoid any ambiguity when taking duals of graded vector spaces, we adopt the conventions outlined in [CDC1, §2.1.1]. By (2.1), we have
@@
When $n=1$, we have $K_{1}(\mathbb{Z}_{S})\otimes\mathbb{Q}=\mathbb{Z}_{S}^{\times}\otimes\mathbb{Q}$, which has dimension $|S|$. When $n\geq 2$, we have $K_{2n-1}(\mathbb{Z}_{S})\otimes\mathbb{Q}=K_{2n-1}(\mathbb{Q})\otimes\mathbb{Q}$, which has dimension 0 for $n$ even and dimension $1$ for $n$ odd [borel].
By the formalism of pro-unipotent groups, it follows that $\mathrm{Lie}(U_{S})$ is a free, negatively graded, pro-nilpotent Lie algebra, which is finite-dimensional in each degree (the grading is negative because of Remark˜2.2). Moreover, the functor $\omega$ induces an equivalence between $\operatorname{MT}(\mathbb{Z}_{S})$ and the category of graded finite-dimensional $\mathbb{Q}$-vector spaces with an action of $\operatorname{Lie}(U_{S})$ compatible with the gradings [delignegoncharov, Prop. 2.2]. By [delignegoncharov, A.15], a system of homogeneous free generators of $\operatorname{Lie}(U_{S})$ is obtained by (non-canonically) lifting a basis of each
@@
Hence, the pro-unipotent group $U_{S}$ is non-canonically isomorphic to the free pro-unipotent group (in the sense of [CDC1, §2.1.3]) on the graded vector space $\bigoplus_{n\in\mathbb{Z}}\operatorname{Ext}^{1}(\mathbb{Q}(0),\mathbb{Q}(n))^{\vee}$ with $\operatorname{Ext}^{1}(\mathbb{Q}(0),\mathbb{Q}(n))^{\vee}$ placed in degree n-n.
@@
Fix once and for all a system of free generators of $\operatorname{Lie}(U_{S})$ given by $\{\tau_{\ell}\}_{\ell\in S}$ in degree (i.e., half-weight) 1-1, and $\sigma_{2n+1}$ in degree 2n1-2n-1, for each $n\geq 1$. We denote by $\Sigma$ this ordered set of free generators, and by $\Sigma_{-d}$ the subset of generators of degree d-d.  [^4] These generators may be chosen arbitrarily, but for the arithmetic step one should choose them as in [CDC1, Prop. 4.1].
@@ 2.3. The unipotent de Rham fundamental group
Fix a point $x\in\mathbb{Q}\setminus\{0,1\}$ or a tangential vector at $0,1$ or $\infty$. The unipotent de Rham fundamental group of $X$ at $x$, $\pi^{\mathrm{un},\operatorname{dR}}_{1}(X;x)$, is the Tannakian fundamental group of the neutral $\mathbb{Q}$-Tannakian category of algebraic vector bundles with nilpotent connection on $X_{\mathbb{Q}}$, with fibre functor given by taking the fibre at $x$. In the case that $x=\vec{1}_{0}$ is the tangential vector $1$ at the point 0, we set
@@
The group $\Pi$ is an affine, pro-unipotent group over $\mathbb{Q}$. Its coordinate ring $\mathcal{O}(\Pi)$ is the de Rham realisation of a commutative Hopf algebra in the category of $\mathrm{ind}$-objects of $\operatorname{MT}(\mathbb{Z})$, the category of unramified mixed Tate motives over $\mathbb{Q}$ [delignegoncharov, §4, §5]. The de Rham realisation functor of $\operatorname{MT}(\mathbb{Z}_{S})$ is canonically isomorphic to $\omega$  [^6] Because the Hodge filtration splits the weight filtration. [delignegoncharov, §2.9]. In particular, the group $\Pi$ carries an action of $G_{S}$.
Let $V$ denote the graded vector space consisting of $H_{1}^{\operatorname{dR}}(X_{\mathbb{Q}})$ in degree 1-1 and 0 in all other degrees. Then we can view $\Pi$ as the free pro-unipotent group on $V$, in the following sense. Take the completed graded tensor algebra $TV$ on $V$ and put the unique coproduct on it such that all elements of $V$ are primitive. Denote by $\mathfrak{n}$ the subspace of $TV$ consisting of primitive elements. Then $\mathfrak{n}$ has the structure of a strictly negatively graded pro-nilpotent Lie algebra, so it is the Lie algebra of a graded pro-unipotent group $U(V)$. The statement that $\Pi$ is the free pro-unipotent group on $V$ means that $\Pi=U(V)$. In particular, the coordinate ring $\mathcal{O}(\Pi)$ is the Hopf dual algebra of the universal enveloping algebra $\mathcal{U}\mathfrak{n}$. This enveloping algebra is $TV$ with the unique coproduct for which all elements of $H_{1}^{\operatorname{dR}}(X_{\mathbb{Q}})$ are primitive.
This description allows us to apply the machinery of [CDC1, §2.1.3]. Take $\{e_{0},e_{1}\}\subseteq TV$ to be the basis dual to the standard basis $\{dz/z,dz/(1-z)\}$ of $H^{1}_{\operatorname{dR}}(X_{\mathbb{Q}})$, assign to these basis elements the degree 1-1 and declare them to be primitive. Then $TV$ is generated by words $w$ in the basis elements $e_{0}$ and $e_{1}$. Let $\operatorname{Li}^{\mathfrak{u}}_{w}$ denote the element of $\mathcal{O}(\Pi)$ that is dual to the word $w$. Then $\mathcal{O}(\Pi)$ is generated by the elements $\{\operatorname{Li}^{\mathfrak{u}}_{w}\}_{w}$. The deconcatenation coproduct is given by
@@ 3.1. The polylogarithmic quotient
Recall from Section˜2.3 that $\Pi$ is the free pro-unipotent group on the graded vector space $V$, which is $H_{1}^{\operatorname{dR}}(X_{\mathbb{Q}})=\mathrm{Span}_{\mathbb{Q}}\{e_{0},e_{1}\}$ in degree 1-1 and zero elsewhere. The specific words of the form
@@ 3.2. The Selmer scheme
[…]
@@ 3.3. The Chabauty–Kim diagram
    $\varphi^{\#}_{b}\colon\mathcal{O}(\Pi)\to\mathbb{Q}_{p},\qquad\operatorname{Li}^{\mathfrak{u}}_{w}\mapsto\int_{b}^{\vec{1}_{0}}\omega_{w},$
@@ 3.4. The Chabauty–Kim locus
Recall from Definition˜3.3 that $\mathcal{I}^{\operatorname{PL}}_{S}=\{f\in\mathcal{O}((\Pi^{\operatorname{PL}})_{K})\mid f(\operatorname{Im}(\operatorname{ev}_{K}))=0\}$. Given an integer $d\geq 1$, we define
@@
Moreover, the set $X(\mathbb{Z}_{p})^{\mathrm{PL}}_{S,d}$ is finite for sufficiently large $d$ (this follows for instance from Proposition˜5.3 below). It is conjectured [CDC1, Conj. 2.32] that
[…]
@@
where $X(\mathbb{Z}_{p})^{\rm Kim}_{S,d}$ denotes Kim’s $p$-adic Chabauty locus in depth $d$ (see [CDC1, Rem. 2.28]). From these inclusions, it is clear that conjecture (3.4) implies Kim’s conjecture [kim]. One insight of [CDC1] is that it may suffice to use single polylogarithms to deal with the $S$-unit equation, as opposed to requiring multiple polylogarithms as in Kim’s original approach. Dan-Cohen and Corwin verified the conjectural equality (3.4) for $S=\{3\}$, $p=5,7$, and $d=4$ [CDC1, Thm. 5.5].
@@
  • Arithmetic step in depth $d$: the task of determining the $p$-adic Chabauty–Kim functions $f|_{X(\mathbb{Z}_{p})}$ defined by (3.2) for the functions $f$ found in the geometric step. This requires the calculation of the image of $f$ under the $p$-adic period map $\operatorname{per}_{p}$. The resulting functions $f|_{X(\mathbb{Z}_{p})}$ can be checked to be non-trivial in practice (and are guaranteed to be non-trivial if one assumes the $p$-adic Period Conjecture [CDC1, Conj. 2.25]). This step depends on the prime $p$ and the specific primes contained in $S$.
@@ 4. The geometric step in coordinates
where we recall that the ideal $\mathcal{I}^{\operatorname{PL}}_{S}=\ker(\operatorname{ev}_{K}^{\#})$ is defined in Definition˜3.3. In this section we describe, following [CDC1], a (non-canonical) affine model for the Selmer scheme $\operatorname{Sel}_{S}^{\operatorname{PL}}$ (Definition˜3.2) and give an explicit formula for the map $\operatorname{ev}$ in the chosen coordinates.
@@ 4.1. Coordinates on Selmer schemes
Recall from Definition˜2.5 the fixed system of free generators $\Sigma=\{\tau_{\ell},\sigma_{2n+1}\}$ of $\operatorname{Lie}(U_{S})$, with their specified degrees. Each word $w$ in these generators corresponds to a dual element in $\mathcal{O}(U_{S})$, denoted $f_{w}$ (the abstract shuffle basis).
@@ 4.2. The evaluation map in coordinates
$\log^{\mathfrak{u}}\mapsto\sum_{\tau\in\Sigma_{-1}}f_{\tau}\Phi_{e_{0}}^{\tau}\qquad\text{ and }\qquad\operatorname{Li}_{k}^{\mathfrak{u}}\mapsto\sum f_{\rho\tau_{(1)}\ldots\tau_{(r)}}\Phi^{\rho}_{e_{1}e_{0}^{s-1}}\Phi_{e_{0}}^{\tau_{(1)}}\ldots\Phi_{e_{0}}^{\tau_{(r)}},$
where the latter sum is taken over $\tau_{(1)},\ldots,\tau_{(r)}\in\Sigma_{-1}$ and $\rho\in\Sigma_{-s}$ such that $r+s=k$ and $1\leq s\leq k$.